OpenStudy (anonymous):

fan and will medal Jo raises purebred wingspaner spaniels. Out of the 37 puppies, he has raised, 6 have gone on to become champion wingspaner spaniels. a. Estimate the odds in favor of Jo’s raising a champion wingspaner spaniel? Show your work. b. Estimate the probability that Jo won’t raise a champion wingspaner spaniel? Show your work.

OpenStudy (anonymous):

a.6/37=0.162162... b.(37-6)/37=31/37=0.837837... Estimate to whatever decimal necessary.

OpenStudy (anonymous):

thankyou